Angle for 3rd maximum when wavelength 580nm and slit seperation 0.1mm
1
Expert's answer
2020-10-15T10:51:04-0400

Solution

According to bragg equation

"2d\sin\theta= n\lambda"

Here given n=3

"\lambda=580 \times 10^{-9}m"

d=0.0001m

So angle of third maximum

"\sin \theta=\frac{3\times580 \times 10^{-9}}{2\times 0.0001}"

"\sin \theta=0.00870"

"\theta=0.50^°"